Solve the following equations:
3p/4 = 6 ….................

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to multiply both sides of the equation by 4,
Then, we get,
= (3p/4) × (4) = 6 × 4
= 3p = 24
Now,
We have to divide both sides of the equation by 3,
Then, we get,
= 3p/3 = 24/3
=p= 8

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve it as follows :
3p / 4 = 6
Cross multiplying, 
3p = 24
P = 24/3
P = 8
 
Hope it helps!
Thank you ✌️